OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

@katragaddasaichandra dont give just answers

OpenStudy (anonymous):

and you gave wrong answer too

OpenStudy (anonymous):

m = (y2-y1)/(x2-x1)=3 y = mx + b b = 5 y = 3x + 7 so its 3x - y = -7

OpenStudy (anonymous):

thank you!